OpenStudy (anonymous):

Write an inequality to represent the real-world problem. Then explain how you created the inequality. Your clothes washer stopped working during the spin cycle and you need to get a person in to fix the washer. Company A costs $20 for the visit and $15 for every hour the person is there to fix the problem. Company B costs $40 for the visit and $5 for every hour the person is there to fix the problem. When would Company B be cheaper than Company A?

OpenStudy (anonymous):

Ok so first start by writing out both equations. Post them here so I can continue helping

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Help someone

OpenStudy (anonymous):

20+ 15x ≤ 40+ 5x the $20 for the visit, $15 for every hour, $40 for the visit, $5 for every hour.

OpenStudy (anonymous):

you have to take the information and rearrange it to get the equation.

OpenStudy (anonymous):

so what you wrote is that the final answer

OpenStudy (anonymous):

I wrote the equations yes.

OpenStudy (anonymous):

k thanks

OpenStudy (anonymous):

your welcome, thank you for working with me now.

OpenStudy (anonymous):

Consider the following statement. Five less than 4 times a number equals 2 more than 3 times the number. a. Write an equation to represent the statement. Use n for the unknown number. b. Solve the equation. c. Explain how you solved the equation.

OpenStudy (anonymous):

help last one

OpenStudy (anonymous):

5-4n=2+3n I think

OpenStudy (anonymous):

u think ?

OpenStudy (anonymous):

plz double check

OpenStudy (anonymous):

I have dyslexia so I might of flipped the less than of grater than

OpenStudy (anonymous):

Five less than 4 times a number equals 2 more than 3 times the number.

OpenStudy (anonymous):

5 > 4x = 2 + 3a

OpenStudy (anonymous):

it said use n for the unknown,

OpenStudy (anonymous):

so whats the final answer

OpenStudy (anonymous):

I have another quesion that states 2 formulas i need to identify them and then ill finish the test

OpenStudy (anonymous):

can you help?

OpenStudy (anonymous):

what are the equations, and how do you need to identify them?

OpenStudy (anonymous):

Problem: Steps: Given problem (1) Identify and discuss the property that was used. (2) Identify and discuss the property that was used

OpenStudy (anonymous):

its 1 problem and the answer is given i just need to identify the property used

OpenStudy (anonymous):

I need to know the problem

OpenStudy (anonymous):

x/2 -5 = 3 x = 16

OpenStudy (anonymous):

idk on this one, sorry

OpenStudy (anonymous):

Thnx anyways and for the problems u helped with srry about before i was extremely stressed

OpenStudy (anonymous):

I understand and know how that is.

OpenStudy (anonymous):

thnx

OpenStudy (anonymous):

no problem, I hope you get better soon.

OpenStudy (anonymous):

thnx can u unblock me?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

just please no swearing at me.

OpenStudy (anonymous):

5-4n=2+3n can you explain how you got this answer

OpenStudy (anonymous):

and no swearing promise

OpenStudy (anonymous):

Five less than 4 times a number equals 2 more than 3 times the number. because the less than would equal the - and grater than would equal the + 5-4n=2+3n

OpenStudy (anonymous):

And Then x/2 - =5-3 which equals x/2 =8 so x/16 after you add like all like terms i just need to know which propertey was used

OpenStudy (anonymous):

for the 1st and seconed part

OpenStudy (anonymous):

the first one is The Commutative Property of Addition. and the second one is The Associative Property of Multiplication.

OpenStudy (anonymous):

how u get that answer

OpenStudy (anonymous):

this jh\Commutative (Order) Property of Addition: Changing the order of the addends does not change the sum. Example: a+b = b+a 3+4 = 4+3 Commutative (Order) Property of Multiplication: Changing the order of the factors does not change the product. Example: ab=ba 5x4 = 4x5

OpenStudy (anonymous):

am I done?

OpenStudy (anonymous):

u good

OpenStudy (anonymous):

ok, good night.

OpenStudy (anonymous):

im done with the thanks allot i will tell you what i get tommoorow good night

OpenStudy (anonymous):

i am done with the test thanks i will tell you what i get tommoroow

OpenStudy (anonymous):

ok, your welcome

OpenStudy (anonymous):

which state do u live in

OpenStudy (anonymous):

Michigan, why?

OpenStudy (anonymous):

what state is that?

OpenStudy (anonymous):

kentucky

OpenStudy (anonymous):

oh, why did you want to know?